caculate all angles of triangle abc if 3times angleb=4 times angle c and angle a =3/7 timesexterior angle a.

Grade:11

2 Answers

Arun
4 years ago
Dear student
 
A = 3(180 - A)/7
 
10A = 3*180
A = 54°
 
Now B +C = 180 - 54 = 126°
 
B = 4C/3
 
Hence 4C/3 + C = 126°
 
7C/3 = 126°
 
C = 54°
 
And B = 4C/3 = 72°
Yashraj
3 years ago
it is a esy question just make equation anduse sum of all angles of triangle is 180 degrees
A = 3(180 - A) divided by 7
A = 54°
we know  B +C = 180 - 54 = 126°
B = 4C/3
so
4C/3 + C = 126°
C = 54°
And B = 4C/3 = 72°